# Modeling isotropic polyconvex hyperelasticity by neural networks – sufficient and necessary criteria for compressible and incompressible materials

### 28 Mar 2026 v1

@@
Abstract: This work investigates different sufficient and necessary criteria for hyperelastic, isotropic polyconvex material models, focusing on neural network implementations for incompressible materials. Furthermore, the expressiveness, accuracy, simplicity as well as the efficiency of those models is analyzed. This also enables an assessment of the practical applicability of the models. Convex Signed Singular Value Neural Networks (CSSV-NNs) are applied to compressible materials and tailored to incompressibility (inc-CSSV-NNs), resulting in a universal approximation for frame-indifferent, isotropic polyconvex energies for the compressible as well as incompressible case. While other existing approaches also guarantee frame-indifference, isotropy and polyconvexity, they impose too restrictive constraints and thus limit the expressiveness of the model since they are only based on sufficient but not necessary criteria. This is further substantiated by numerical examples of several, well-established classical models (Neo–Hooke, Mooney–Rivlin, Gent and Arruda–Boyce) and Treloar’s experimental data. Moreover, the numerical examples include an explicitly constructed energy function that cannot be approximated by neural networks constrained by Ball’s criterion for polyconvexity. This substantiates that Ball’s criterion, though sufficient, is not necessary for polyconvexity.
Keywords: artificial neural networks, hyperelasticity, polyconvexity, incompressibility, constitutive modeling, singular values
@@ 1 Introduction
Due to this observation, Geuken et al. [16] introduced the Convex Signed Singular Value Neural Network (CSSV-NN) architecture, which is based on a necessary and sufficient criterion for frame-indifferent, isotropic polyconvex energies derived by Wiedemann and Peter [45, 46] and can universally approximate any such energy function to arbitrary precision. In contrast, many existing approaches [38, 25, 27, 6, 11, 26, 22, 7, 44, 2] either fail to strictly enforce frame-indifference, isotropy and polyconvexity or do so under overly restrictive sufficient criteria, thereby limiting their expressiveness as stated above.
This work systematically analyzes sufficient and necessary criteria for hyperelastic, frame-indifferent, isotropic and polyconvex energies and investigates their implications for neural network models. This also enables an assessment of the practical applicability of the models. The CSSV-NNs are further extended and adapted to the incompressible case (denoted by inc-CSSV-NN). This is particularly relevant for synthetic materials such as polymers [43, 5, 41] and biological materials such as the periodontal ligament [37, 24]. As for the compressible case, existing formulations of frame-indifferent, isotropic polyconvex energies for incompressibility (e.g., [39, 49, 18, 14, 21, 19]) are too restrictive or do not strictly enforce frame-indifference, isotropy and polyconvexity. This is likely because most of these approaches are primarily designed for inelastic material behavior.
@@
  • Extension and adaptation of Convex Signed Singular Value Neural Networks (CSSV-NNs) to incompressible materials (inc-CSSV-NNs) and other neural network architectures built upon the investigated criteria.
@@
The paper is organized as follows. Section 2 constitutes physical and mathematical constraints for hyperelastic modeling and points out how to achieve frame-indifference, isotropy and incompressibility. Based on that, four different criteria for frame-indifferent, isotropic polyconvex energies of compressible and incompressible materials are presented in Section 3. Those are then used for designing appropriate neural networks for incompressible material behavior in Section 4. Section 5 provides a proof of the universal approximation theorem of inc-CSSV-NNs for frame-indifferent, isotropic polyconvex functions of incompressible materials. The algorithmic implementation in terms of network activation, hyperparameters and training is described in Section 6. Finally, numerical examples of several, well-established classical models and experimental data, such as Treloar’s data, are discussed in Section 7. Moreover, the numerical examples include an explicitly constructed energy function that cannot be approximated by neural networks constrained by Ball’s criterion for polyconvexity.
@@ 2 Physical and mathematical constraints for hyperelastic modeling
The formulation of hyperelastic material models should respect fundamental physical principles and mathematical restrictions. In this work, the focus is on compressible and incompressible, isotropic hyperelastic materials.
@@ 2.1 The compressible case
    Thermodynamic Consistency: The second law of thermodynamics must hold (which is automatically satisfied in hyperelasticity if stresses derive from $\Psi$):
@@
    $\Psi(\boldsymbol{F})=\Psi(\boldsymbol{Q}\cdot\boldsymbol{F})\quad\forall\,\boldsymbol{Q}\in\mathcal{SO}(3).$ (3)
@@
    $\boldsymbol{P}\cdot\boldsymbol{F}^{\mathsf{T}}=2\boldsymbol{F}\cdot\frac{\partial\Psi}{\partial\boldsymbol{C}}\cdot\boldsymbol{F}^{\mathsf{T}}=\boldsymbol{F}\cdot\boldsymbol{P}^{\mathsf{T}},$ (4)
@@
which guarantees the existence of a minimizer in the set of admissible displacements $\boldsymbol{u}$. Foundational work by Morrey [32, 33] and later extensions by Meyers [30], Acerbi and Fusco [3] and Marcellini [29] established that $I$ is lower semicontinuous if $\Psi$ is quasiconvex and satisfies suitable growth and coercivity conditions (see [13, Theorem 8.29]). However, these conditions exclude energies that take the value $\infty$, making them incompatible with (6a) and (6b). To overcome this, Ball [8, 10] introduced polyconvexity as a weaker convexity notion that still guarantees lower semicontinuity and existence of minimizers under milder assumptions that are in particular compatible with the growth conditions (6a) and (6b) (see also [13, Theorem 8.31]). While polyconvexity implies quasiconvexity and rank-one convexity, it is easier to verify and implement. Moreover, it accommodates all physical constraints, making it particularly suitable for modeling elasticity.
@@
2.1.1 Frame-indifference and isotropy
Frame-indifference can be guaranteed a priori by formulating the energy in terms of signed singular values $\nu_{i}$ or principal stretches $\lambda_{i}$, since the singular values of $\boldsymbol{F}$ are equal to the principal stretches and thus the square root of the frame-indifferent eigenvalues of $\boldsymbol{C}$. They can become negative by rotation or – to be more precise – because the decomposition
of $\boldsymbol{F}$ into two rotations $\boldsymbol{R}_{1}$, $\boldsymbol{R}_{2}\in\mathcal{SO}(3)$ and a diagonal matrix $\textbf{diag}(\boldsymbol{\nu})$ is only unique up to permutations and an even number of sign changes for the entries $\nu_{1},\nu_{2},\nu_{3}$ as also motivated by Fig. 1. However, the choice of their sign does not affect the determinant (orientation preserving), yielding $\nu_{1}\,\nu_{2}\,\nu_{3}=\det\boldsymbol{F}=J>0$.
For every frame-indifferent and isotropic function $\Psi(\boldsymbol{F})$ there exists a unique function $\widetilde{\Psi}(\nu_{1},\nu_{2},\nu_{3})$ that characterizes $\Psi$ in terms of the signed singular values of $\boldsymbol{F}$.
This implies that $\widetilde{\Psi}$ is $\Pi_{3}$-invariant with
Analogously, for every frame-indifferent and isotropic function $\Psi(\boldsymbol{F})$ there exists a unique function $\breve{\Psi}(\lambda_{1},\lambda_{2},\lambda_{3})$ that characterizes $\Psi$ in terms of the principal stretches (singular values) of $\boldsymbol{F}$.
@@ 2.2 The incompressible case
to formulate $\Psi^{\textnormal{inc}}$ only in terms of $\boldsymbol{F}^{\textnormal{iso}}$ and thus directly account for the incompressibility constraint [15]. In this work, this split is omitted and $\Psi^{\textnormal{inc}}$ is evaluated only for deformation states $\boldsymbol{F}=\boldsymbol{F}^{\textnormal{iso}}$ ($J=\det\boldsymbol{F}=1$) as, e.g., in [9]. This applies to all sections and examples concerning incompressibility.
To apply the concept of polyconvexity, $\Psi^{\textnormal{inc}}$ is identified with its extension by $\infty$, which is given by $\Psi^{\textnormal{inc}}_{\infty}\colon\mathbb{R}^{3\times 3}\to\mathbb{R}_{\infty}$
@@
For the existence of a minimizer of the energy functional $I$ in the incompressible case, one is interested in $\Psi^{\textnormal{inc}}_{\infty}$ to be polyconvex. This motivates the following definition: A function $\Psi^{\textnormal{inc}}$ defined on the subset $\{\boldsymbol{F}\in\mathbb{R}^{3\times}\mid\det(\boldsymbol{F})=1\}$ is called polyconvex if the corresponding function $\Psi^{\textnormal{inc}}_{\infty}$ given by Eq. (15) is polyconvex. This extension resembles the infinite stiffness of incompressible materials against volumetric deformations. While the extension is not of much practical relevance, it is an important prerequisite to prove the necessity of the polyconvexity criterion based on signed singular values in Section 3.1.2.
@@
Figure 2: Overview and relations between the investigated criteria for frame-indifferent, isotropic polyconvex energies of compressible materials. For incompressibility the determinant’s influence via arguments $\nu_{1}\nu_{2}\nu_{3},\lambda_{1}\lambda_{2}\lambda_{3}$ and $I_{3}^{\boldsymbol{U}}$ is omitted. The relationships regarding the sufficiency and necessity are the same in the incompressible case here, which is not generally the case. The ?-marked relation is indicated by the numerical results in Section 7.
@@ 3.1.1 The compressible case
Then, Theorem 1.4 in [46] provides a sufficient and necessary criterion for frame-indifferent, isotropic polyconvex energies:
@@
The proof of Theorem 1 was given by Wiedemann and Peter [45, 46]. The first three arguments of $\Psi_{\textnormal{ssv}}$ are connected to the convexity in $\boldsymbol{F}$, the next three to the one in $\textnormal{cof}\,\boldsymbol{F}$ and the last argument is the determinant $\textnormal{det}\,\boldsymbol{F}$. Using
@@ 3.1.2 The incompressible case
In this section, incompressible energies $\Psi^{\textnormal{inc}}$ are considered, as described in Section 2.2. With [46, Lemma 2.4] one can identify incompressible functions $\Psi^{\textnormal{inc}}$ that are defined on $\{\boldsymbol{F}\in\mathbb{R}^{3\times 3}\mid J=\det(\boldsymbol{F})=1\}$ with $\Pi_{3}$-invariant functions $\widetilde{\Psi}^{\textnormal{inc}}$ defined on $\{\boldsymbol{\nu}\in\mathbb{R}^{3}\mid\nu_{1}\nu_{2}\nu_{3}=1\}$. In line with the compressible case, a function $\widetilde{\Psi}^{\textnormal{inc}}$ defined on $\boldsymbol{\nu}\in\mathbb{R}^{3}$ with $\nu_{1}\nu_{2}\nu_{3}=1$ is called singular value polyconvex if the corresponding function $\Psi^{\textnormal{inc}}$ is polyconvex. Recap that the polyconvexity of $\Psi^{\textnormal{inc}}$ is defined by the polyconvexity of $\Psi^{\textnormal{inc}}_{\infty}$. Thus, $\widetilde{\Psi}^{\textnormal{inc}}$ is singular value polyconvex if and only if its extension by $\infty$, which is given by $\widetilde{\Psi}^{\textnormal{inc}}_{\infty}\colon\mathbb{R}^{3}\to\mathbb{R}_{\infty}$
@@
For the other direction, let $\Psi_{\textnormal{ssv}}^{\textnormal{inc}}\colon\mathbb{R}^{6}\to\mathbb{R}_{\infty}$ be a convex and lower semicontinuous function that satisfies Eq. (22). The idea is to construct a convex and lower semicontinuous function $\Psi_{\textnormal{ssv}}\colon\mathbb{R}^{7}\rightarrow\mathbb{R}_{\infty}$ that satisfies Eq. (18) for $\widetilde{\Psi}=\widetilde{\Psi}^{\textnormal{inc}}_{\infty}$. Then, the singular value polyconvexity of $\widetilde{\Psi}^{\textnormal{inc}}$ follows from Theorem 1. Define $\Psi_{\textnormal{ssv}}$ by $\Psi_{\textnormal{ssv}}(\boldsymbol{x})\coloneqq\Psi_{\textnormal{ssv}}^{\textnormal{inc}}(\boldsymbol{x})$ for $x_{7}=1$ and $\Psi_{\textnormal{ssv}}(\boldsymbol{x})\coloneqq\infty$ for $x_{7}\neq 1$. This function is the extension of a convex function outside the convex domain $\{\boldsymbol{x}\in\mathbb{R}^{7}\mid x_{7}=1\}$ by $\infty$ and, thus, it is convex. Since $\Psi_{\textnormal{ssv}}^{\textnormal{inc}}$ is lower semicontinuous and the set $\{\boldsymbol{x}\in\mathbb{R}^{7}\mid x_{7}=1\}$ is closed, the extension by $\infty$ is lower semicontinuous. It remains to check Eq. (18) for $\widetilde{\Psi}=\widetilde{\Psi}^{\textnormal{inc}}_{\infty}$:

### 13 Apr 2026 v2

@@
Abstract: This work investigates different sufficient and necessary criteria for hyperelastic, isotropic polyconvex material models, focusing on neural network implementations for compressible and incompressible materials. Furthermore, the expressiveness, accuracy, simplicity as well as the efficiency of those models is analyzed. This also enables an assessment of the practical applicability of the models. Convex Signed Singular Value Neural Networks (CSSV-NNs) are applied to compressible materials and tailored to incompressibility (inc-CSSV-NNs), resulting in a universal approximation for frame-indifferent, isotropic polyconvex energies for the compressible as well as incompressible case. While other existing approaches also guarantee frame-indifference, isotropy and polyconvexity, they impose too restrictive constraints and thus limit the expressiveness of the model. This is further substantiated by numerical examples of several, well-established classical models (Neo–Hooke, Mooney–Rivlin, Gent and Arruda–Boyce) and Treloar’s experimental data. Moreover, the numerical examples include an explicitly constructed energy function that cannot be approximated by neural networks constrained by an improved version of Ball’s criterion for polyconvexity. This substantiates that Ball’s criterion, though sufficient, is not necessary for polyconvexity.
Keywords: artificial neural networks, hyperelasticity, polyconvexity, incompressibility, singular values, constitutive modeling
@@ 1 Introduction
Due to this observation, Geuken et al. [16] introduced the Convex Signed Singular Value Neural Network (CSSV-NN) architecture, which is based on a necessary and sufficient criterion for frame-indifferent, isotropic polyconvex energies derived by Wiedemann and Peter [45, 46]. They can universally approximate any such energy function to arbitrary precision. In contrast, many existing approaches [38, 25, 27, 6, 11, 26, 22, 7, 44, 2] either fail to strictly enforce frame-indifference, isotropy and polyconvexity or do so under overly restrictive sufficient criteria, thereby limiting their expressiveness as stated above.
This work systematically analyzes sufficient and necessary criteria for hyperelastic, frame-indifferent, isotropic and polyconvex energies and investigates their implications for neural network models. This also enables an assessment of the practical applicability of the models. The CSSV-NNs are further extended and adapted to the incompressible case (denoted by inc-CSSV-NN). This is particularly relevant for synthetic materials such as polymers [43, 5, 41] and biological materials such as the periodontal ligament [37, 24]. As for the compressible case, existing formulations of frame-indifferent, isotropic polyconvex energies for incompressibility (e.g., [39, 49, 18, 14, 21, 19]) are too restrictive or do not strictly enforce frame-indifference, isotropy and polyconvexity.
@@
  • Extension and adaptation of Convex Signed Singular Value Neural Networks (CSSV-NNs) to incompressible materials (inc-CSSV-NNs).
@@
The paper is organized as follows. Section 2 constitutes physical and mathematical constraints for hyperelastic modeling and points out how to achieve frame-indifference, isotropy and incompressibility. Based on that, four different criteria for frame-indifferent, isotropic polyconvex energies of compressible and incompressible materials are presented in Section 3. Those are then used for designing appropriate neural networks for incompressible material behavior in Section 4. Section 5 provides a proof of the universal approximation theorem of inc-CSSV-NNs for frame-indifferent, isotropic polyconvex functions of incompressible materials. The algorithmic implementation in terms of network activation, hyperparameters and training is described in Section 6. Finally, numerical examples of several, well-established classical models and experimental data, such as Treloar’s data, are discussed in Section 7. Moreover, the numerical examples include an explicitly constructed energy function that cannot be approximated by neural networks constrained by an improved version of Ball’s criterion for polyconvexity.
@@ 2 Physical and mathematical constraints for hyperelastic modeling
Hyperelastic material models should respect fundamental physical principles and mathematical restrictions. In this work, the focus is on compressible and incompressible, isotropic hyperelastic materials.
@@ 2.1 The compressible case
    Thermodynamic Consistency: The second law of thermodynamics must hold (which is automatically satisfied in hyperelasticity if stresses derive from $\Psi$), reading here:
@@
    $\Psi(\boldsymbol{F})$ $=\Psi(\boldsymbol{Q}\cdot\boldsymbol{F})\quad\forall\,\boldsymbol{Q}\in\mathcal{SO}(3),$ (3a)
@@
    $\boldsymbol{P}\cdot\boldsymbol{F}^{\mathsf{T}}$ $=2\boldsymbol{F}\cdot\frac{\partial\Psi}{\partial\boldsymbol{C}}\cdot\boldsymbol{F}^{\mathsf{T}}=\boldsymbol{F}\cdot\boldsymbol{P}^{\mathsf{T}},$ (3b)
@@
which guarantees the existence of a minimizer in the set of admissible displacements $\boldsymbol{u}$. Foundational work by Morrey [32, 33] and later extensions by Meyers [30], Acerbi and Fusco [3] and Marcellini [29] established that $I$ is lower semicontinuous if $\Psi$ is quasiconvex and satisfies suitable growth and coercivity conditions (see [13, Theorem 8.29]). However, these conditions exclude energies that take the value $\infty$, making them incompatible with (5a) and (5b). To overcome this, Ball [8, 10] introduced polyconvexity as a weaker convexity notion that still guarantees lower semicontinuity and existence of minimizers under milder assumptions that are in particular compatible with the growth conditions (5a) and (5b) (see also [13, Theorem 8.31]). While polyconvexity implies quasiconvexity and rank-one convexity, polyconvexity is easier to verify and implement. Moreover, it accommodates all physical constraints, making it particularly suitable for modeling elasticity.
@@ 2.2 The incompressible case
to formulate $\Psi^{\textnormal{inc}}$ only in terms of $\boldsymbol{F}^{\textnormal{iso}}$ and thus directly account for the incompressibility constraint [15]. In this work, this split is omitted and $\Psi^{\textnormal{inc}}$ is evaluated only for deformation states $\boldsymbol{F}=\boldsymbol{F}^{\textnormal{iso}}$ ($J=\det\boldsymbol{F}=1$) as, e.g., in [9]. This holds for all sections and examples concerning incompressibility.
To apply the concept of polyconvexity, $\Psi^{\textnormal{inc}}$ is identified with its extension by $\infty$, which is given by $\Psi^{\textnormal{inc}}_{\infty}\colon\mathbb{R}^{3\times 3}\to\mathbb{R}_{\infty}$ as
@@
For the existence of a minimizer of the energy functional $I$ in the incompressible case, one is interested in $\Psi^{\textnormal{inc}}_{\infty}$ to be polyconvex. This motivates the following definition:
A function $\Psi^{\textnormal{inc}}$ defined on the subset $\{\boldsymbol{F}\in\mathbb{R}^{3\times}\mid\det(\boldsymbol{F})=1\}$ is called polyconvex if the corresponding function $\Psi^{\textnormal{inc}}_{\infty}$ given by Eq. (11) is polyconvex.
This extension resembles the infinite stiffness of incompressible materials against volumetric deformations. The practical relevance of this extension is two-fold. First, it allows to apply polyconvexity criteria to the incompressible case with minor modifications of the criteria themselves. Second, it is an important prerequisite to prove the necessity of the polyconvexity criterion based on signed singular values in Section 3.1.2.
@@
2.3 Energy parametrizations in principal stretches and signed singular values
Frame-indifference can be guaranteed a priori by formulating the energy in terms of different arguments, e.g., in terms of eigenvalues of $\boldsymbol{C}$, principal stretches $\lambda_{i}$ or signed singular values $\nu_{i}$ of $\boldsymbol{F}$. Starting with the eigenvalues of $\boldsymbol{C}$, the principal stretches are obtained as their square roots, $\lambda_{i}=\sqrt{\operatorname{eig}(\boldsymbol{C})_{i}}$. The principal stretches are moreover equal to the singular values of the deformation gradient $\boldsymbol{F}$. Eventually, the principal stretches (or singular values) are the absolute values of the signed singular values of $\boldsymbol{F}$, $\lambda_{i}=|\nu_{i}|$. Note that the signed singular values $\nu_{i}$ contain most (physical) information, which is lost for the other arguments by transformation into absolute or quadratic values.
In contrast to the stretches, the signed singular values can become negative by rotation or – to be more precise – because the decomposition
of $\boldsymbol{F}$ into two rotations $\boldsymbol{R}_{1}$, $\boldsymbol{R}_{2}\in\mathcal{SO}(3)$ and a diagonal matrix $\textbf{diag}(\boldsymbol{\nu})$ is only unique up to permutations and an even number of sign changes for the entries $\nu_{1},\nu_{2},\nu_{3}$ as also motivated by Fig. 1. However, the choice of their sign does not affect the determinant (orientation preserving), yielding $\nu_{1}\,\nu_{2}\,\nu_{3}=\det\boldsymbol{F}=J>0$. Signed singular values will be of particular importance for the following investigation. A key property for this is:
For every frame-indifferent and isotropic function $\Psi(\boldsymbol{F})$ there exists a unique function $\widetilde{\Psi}(\nu_{1},\nu_{2},\nu_{3})$ that characterizes $\Psi$ in terms of the signed singular values of $\boldsymbol{F}$, cf. [46].
This implies that $\widetilde{\Psi}$ is $\Pi_{3}$-invariant with
Analogously, for every frame-indifferent and isotropic function $\Psi(\boldsymbol{F})$ there exists a unique function $\breve{\Psi}(\lambda_{1},\lambda_{2},\lambda_{3})$ that characterizes $\Psi$ in terms of the principal stretches (singular values) of $\boldsymbol{F}$, cf. [46].
@@ 2.3 Energy parametrizations in principal stretches and signed singular values
Figure 2: Overview and relations between the investigated criteria for frame-indifferent, isotropic polyconvex energies of compressible materials. For incompressibility the determinant’s influence via arguments $\nu_{1}\nu_{2}\nu_{3},\lambda_{1}\lambda_{2}\lambda_{3}$ and $I_{3}^{\boldsymbol{U}}$ is omitted. The open, ?-marked relation is indicated by the numerical results in Section 7.
@@ 3.1.1 The compressible case
Then, Theorem 1.4 in [46] provides a sufficient and necessary criterion for frame-indifferent, isotropic polyconvex energies as follows:
@@
The proof of Theorem 1 was given by Wiedemann and Peter [45, 46]. The first three arguments of $\Psi_{\textnormal{ssv}}$ are connected to the convexity with respect to $\boldsymbol{F}$, the next three to the convexity with respect to $\textnormal{cof}\,\boldsymbol{F}$ and the last argument is the determinant $\textnormal{det}\,\boldsymbol{F}$. Using
@@ 3.1.2 The incompressible case
In this section, incompressible energies $\Psi^{\textnormal{inc}}$ are considered, as described in Section 2.2. With [46, Lemma 2.4] one can identify incompressible functions $\Psi^{\textnormal{inc}}$ that are defined on $\{\boldsymbol{F}\in\mathbb{R}^{3\times 3}\mid J=\det(\boldsymbol{F})=1\}$ with $\Pi_{3}$-invariant functions $\widetilde{\Psi}^{\textnormal{inc}}$ defined on $\{\boldsymbol{\nu}\in\mathbb{R}^{3}\mid\nu_{1}\nu_{2}\nu_{3}=1\}$. In line with the compressible case, a function $\widetilde{\Psi}^{\textnormal{inc}}$ defined on $\boldsymbol{\nu}\in\mathbb{R}^{3}$ with $\nu_{1}\nu_{2}\nu_{3}=1$ is called singular value polyconvex if the corresponding function $\Psi^{\textnormal{inc}}$ is polyconvex. Recall that the polyconvexity of $\Psi^{\textnormal{inc}}$ is defined by the polyconvexity of $\Psi^{\textnormal{inc}}_{\infty}$. Thus, $\widetilde{\Psi}^{\textnormal{inc}}$ is singular value polyconvex if and only if its extension by $\infty$, which is given by $\widetilde{\Psi}^{\textnormal{inc}}_{\infty}\colon\mathbb{R}^{3}\to\mathbb{R}_{\infty}$
@@
For the other direction, let $\Psi_{\textnormal{ssv}}^{\textnormal{inc}}\colon\mathbb{R}^{6}\to\mathbb{R}_{\infty}$ be a convex and lower semicontinuous function that satisfies Eq. (21). The idea is to construct a convex and lower semicontinuous function $\Psi_{\textnormal{ssv}}\colon\mathbb{R}^{7}\rightarrow\mathbb{R}_{\infty}$ that satisfies Eq. (17) for $\widetilde{\Psi}=\widetilde{\Psi}^{\textnormal{inc}}_{\infty}$. Then, the singular value polyconvexity of $\widetilde{\Psi}^{\textnormal{inc}}$ follows from Theorem 1. Define $\Psi_{\textnormal{ssv}}$ by $\Psi_{\textnormal{ssv}}(\boldsymbol{x})\coloneqq\Psi_{\textnormal{ssv}}^{\textnormal{inc}}(\boldsymbol{x})$ for $x_{7}=1$ and $\Psi_{\textnormal{ssv}}(\boldsymbol{x})\coloneqq\infty$ for $x_{7}\neq 1$. This function is the extension of a convex function outside the convex domain $\{\boldsymbol{x}\in\mathbb{R}^{7}\mid x_{7}=1\}$ by $\infty$ and, thus, it is convex. Since $\Psi_{\textnormal{ssv}}^{\textnormal{inc}}$ is lower semicontinuous and the set $\{\boldsymbol{x}\in\mathbb{R}^{7}\mid x_{7}=1\}$ is closed, the extension by $\infty$ is lower semicontinuous. It remains to check Eq. (17) for $\widetilde{\Psi}=\widetilde{\Psi}^{\textnormal{inc}}_{\infty}$. It holds, since: